OpenStudy (mathisfun13):

Find the distance from point A (3,-2) to point B (7,1)

ganeshie8 (ganeshie8):

distance formula

ganeshie8 (ganeshie8):

distance = \(\sqrt{(x2-x1)^2 + (y2-y1)^2}\)

ganeshie8 (ganeshie8):

A (3,-2) to point B (7, 1) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plugin

OpenStudy (mathisfun13):

okay what about (3√7, 2√5) to point B (5√7, 5√5)

ganeshie8 (ganeshie8):

same formula

ganeshie8 (ganeshie8):

dont get distracted wid the radical signs. they will go away once u take square

OpenStudy (mathisfun13):

okay

ganeshie8 (ganeshie8):

start by plugging them in the formula

OpenStudy (mathisfun13):

i got (2√7) +(3 √5)

ganeshie8 (ganeshie8):

thats not all... keep going...

OpenStudy (mathisfun13):

how do u square that tho?

ganeshie8 (ganeshie8):

\(\sqrt{ (2√7)^2 +(3 √5)^2}\)

OpenStudy (mathisfun13):

do i square only the 7 and the 5?

ganeshie8 (ganeshie8):

\((2\sqrt{7})^2\)

ganeshie8 (ganeshie8):

square applies to both \(2\) and \(\sqrt{7}\)

ganeshie8 (ganeshie8):

\((2\sqrt{7})^2 = 2^2\sqrt{7} ^2\)

OpenStudy (mathisfun13):

4√49+9√25

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

but u can simplify more

ganeshie8 (ganeshie8):

\(\sqrt{49} = ?\)

OpenStudy (mathisfun13):

7

ganeshie8 (ganeshie8):

\(\sqrt{25} = ?\)

OpenStudy (mathisfun13):

5

ganeshie8 (ganeshie8):

yes

OpenStudy (mathisfun13):

how do i put that in decimal form rounded to the nearest hundredth?

ganeshie8 (ganeshie8):

\(\sqrt{ (2√7)^2 +(3 √5)^2}\) \(\sqrt{ 4 (7) +9(5)}\)

ganeshie8 (ganeshie8):

\(\sqrt{ (2√7)^2 +(3 √5)^2}\) \(\sqrt{ 4 (7) +9(5)}\) \(\sqrt{ 28 +45}\)

ganeshie8 (ganeshie8):

\(\sqrt{ (2√7)^2 +(3 √5)^2}\) \(\sqrt{ 4 (7) +9(5)}\) \(\sqrt{ 28 +45}\) \(\sqrt{73}\)

ganeshie8 (ganeshie8):

use ur calculator and find \(\sqrt{73}\)

OpenStudy (mathisfun13):

i got 73

ganeshie8 (ganeshie8):

yes, take square root of it

OpenStudy (mathisfun13):

so √73?

ganeshie8 (ganeshie8):

yes, do u have calculator ?

OpenStudy (mathisfun13):

yess

OpenStudy (mathisfun13):

thats what it showed me. so then .73?

ganeshie8 (ganeshie8):

good :) take square root of 73 and see wat u get

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

On ur calculator keys, you should see a symbol like that...

ganeshie8 (ganeshie8):

thats the symbol for square root.

OpenStudy (mathisfun13):

okay. i got 8.544003745317351

ganeshie8 (ganeshie8):

how do i put that in decimal form rounded to the nearest hundredth?

ganeshie8 (ganeshie8):

thats the decimal form. now you still need to round it to hundredth

ganeshie8 (ganeshie8):

8.544003745317351 ^

ganeshie8 (ganeshie8):

rounding to hundredth makes it : 8.54